# Incommensurate and commensurate antiferromagnetic orders in the kagome compound UV6Sn6

### v1

@@ Abstract
We report on the synthesis of single crystals of a new kagome compound, UV6Sn6, and present the results of magnetization, electrical resistivity, heat capacity, X-ray and neutron diffraction experiments to characterize the structure and magnetic properties. UV6Sn6 crystallizes in a large supercell of the HfFe6Ge6 parent structure with an hexagonal symmetry in which some of the U atoms are shifted by $c/2$ in an ordered fashion. Below $T_{N1}=30$ K, an incommensurate magnetic structure with a temperature-dependent wavevector $(0,0,k_{z})$ is observed. Below $T_{N2}=23.5$ K, the wavevector locks in to $(0,0,0.5)$ forming an antiferromagnetic ground state. The U moments align along the $c$ axis retaining a large magnetic anisotropy.
@@ I Introduction
Materials containing kagome networks have attracted a lot of attention for the study of topologically nontrivial phases, and intertwined orders [1, 2, 3]. Among them, the 166 family of compounds, characterized by the general formula $RT_{6}X_{6}$, represents a fertile platform for investigating the rich interplay of topology, electron correlations, and magnetism. The R element can be a rare-earth element or Li, Mg, Ca, Ti, Zr, Hf, T is a transition metal element, and X is typically a tetrel element such as Sn or Ge. In these materials, the kagome lattice, a two-dimensional network of corner-sharing triangles, facilitates the emergence of exotic phenomena, including charge density waves [4, 5, 6], phonon softening [7, 8, 9], gapped Dirac fermions and surface states [10, 11, 12], a nodal ring [13], anomalous and topological Hall effects [14, 15, 16], anomalous Nernst effect [17, 18], and an extraordinary structural diversity [19, 20, 21, 22, 23, 24, 25].
@@ II Methods
Single crystals of UV6Sn6 were grown using a self-flux method similar to that reported for YV6Sn6 [11]. Uranium (cleaned by electrolytic etching to remove the oxide [26]), vanadium pieces, tin shots (cleaned with 30 vol.HCl) were introduced in a Canfield crucible set [27] in the 1:6:20 molar ratio, and sealed in a silica ampoule with 1/3 atm argon gas. The ampoule was heated to 1150 °C in 7 hours, held there for 36 hours, slowly cooled down to 1050 °C in 287 hours, after which it was decanted using a centrifuge. Only crystals of UV6Sn6 were obtained with this temperature profile. The experiments with cooling to 780 °C resulted in additional single crystals of USn3.
@@
Single-crystal X-ray diffraction data were collected at 100–150 K using the four-circle Kappa geometry diffractometers: Rigaku XCallibur (Rigaku CrysAlis CCD software), and Bruker D8 Venture (APEX software). In both instruments a Mo $\text{Mo }K_{\alpha}$ radiation source was used (0.71073 Å). The structural refinement was carried out using the crystallography software package GSAS-II [28].
@@
Magnetization measurements were performed using an MPMS3 SQUID magnetometer (Quantum Design) in VSM scan mode. The crystallographic axes of the sample were determined using Laue diffraction followed by polishing to obtain a flat surface perpendicular to the $c$-axis.
@@
Resistivity measurements were performed in a physical property measurement system PPMS (Quantum Design) using a usual four point lock-in technique; 15 $\mu$m Au wires have been spot welded to the sample for electrical contacts. Due to the irregular shape of the sample and the not perfect alignment of the electrical contacts, absolute values of the resistivity were not determined. The specific heat has been measured using the PPMS with heat pulse relaxation method; heat pulses of 0.5, 1, and 2% of the actual temperature gave similar results. The samples for specific heat had masses of 1.45 and 0.9 mg, thus the specific heat of the sample is small compared to the addenda contribution.
@@
Neutron diffraction experiments were carried out at Institut Laue Langevin, Grenoble on the two axis diffractometer D23 and the large solid angle detector diffractometer XtremeD. A single-crystal sample with a mass of 1.48 mg was pre-aligned using X-ray diffraction (Laue method) with the $b$-axis oriented vertically so that the horizontal scattering plane was ($a^{*}$, $c^{*}$) in hexagonal setting. D23 is equipped with a lifting detector reaching 30° and neutrons of wavelength 1.28 Å and 2.36 Å were used with a copper monochromator and a pyrolitic graphite (PG) monochromator, respectively. In this latter case, a PG filter was inserted in the incoming beam in order to remove higher order contamination. On XtremeD, the wavelength of 2.45 Å provided by a PG monochromator was used, the angular coverage of the detector is of 130° in the horizontal plane and 24° in the vertical plane. A PG filter was also installed. An oscillating radial collimator is placed in front of the detector bank in order to suppress the sample environment contribution to the collected data and reduce background. For both experiments the sample was inserted in an helium flow cryostat. The Mag2POL software was used to refine the nuclear and magnetic structures [29].
@@ III.1 Crystal Structure
Figure 1: The HfFe6Ge6-type parent structure of UV6Sn6, where the V atoms (red) form a kagome network (a), and the U atoms (blue) form a triangular lattice (b). The V kagome and U triangular lattices arrange in layers to form the structure depicted in (c). The V and Sn atoms around U form a hexagonal cavity represented as a polyhedron in (c). The ordering of these cavities can lead to patterns such as the one shown in (d). These patterns extend the structures of $RT_{6}X_{6}$ compounds into supercells, like the one depicted in (e). The structure of UV6Sn6 can be described as a supercell where a periodic modulation of the HfFe6Ge6-type parent cell—marked in black in (e), yields a 7 × 7 superstructure on the $ab$ plane. The hexagonal cavities occupied by U shift along the $c$ axis by $c/2$ yielding the pattern shown in (d). This pattern is identical along the $b$ crystallographic axis ($a=b$).
@@
Although the goodness-of-fit value of 1.76 (GOF, Table 1) suggests a good agreement between the measured and calculated average intensities, the high value of the weighted profile factor, $wR$ = 20.76%, reveals a significant discrepancy between the observed and calculated structure factors (details in Appendix B).
@@
Careful examination of both the X-ray and neutron diffraction data revealed that the discrepancies with the HfFe6Ge6-type model arise from very weak intensity at some families of Bragg reflections and the presence of additional satellite reflections around them. Notably, it was at the instrument XtremeD at ILL that we first observed these satellite reflections (see Appendix C), motivating further x-ray diffraction experiments. The satellite reflections are particularly evident in the reconstruction of reciprocal space (precession images calculated from diffraction data) centered on planes $hkl$ with $l$ = 3 and $h$ = 2, from X-ray diffraction data (Fig. 2b,e). On the $hk3$ plane (Fig. 2b), we clearly observe groups of satellite reflections around the expected Bragg reflections (e.g. 003, 103, and 013 marked in blue in Fig. 2a), as well as along the $h$ and $k$ directions, forming a triangular grid. Meanwhile, on the $2kl$ plane (Fig. 2e), the apparent discrepancy in the expected peak intensity and the presence of satellite reflections are evident for all odd values of $l$ (i.e., $l$ = 1, 3, 5, etc.).
@@
Figure 2: Comparison of the calculated and observed reconstructions of the reciprocal space centered on planes $hk3$ and $2kl$ for UV6Sn6. The calculated diffraction patterns in (a) and (d) show the expected Bragg spots (reflections, marked in blue) for a HfFe6Ge6-type cell on planes $hk3$ and $2kl$, respectively. The observed $hk3$ pattern in (b) shows undetectable intensity at the expected Bragg spots and the presence of six satellite reflections around them. Additional satellite reflections are observed along $h$ and $k$ directions. The observed $2kl$ pattern in (e) shows discrepancies in the intensity of the expected Bragg reflections, as well as satellite peaks on either side of the Bragg spots along ky. These unexpected features are true for every $l$ = odd value (i.e., $l$ = 1, 3, 5, etc.). Our simulated diffraction patterns for an expanded $7\times 7$ cell (Fig. 1d-e) accurately replicate the observed diffracted spots on the $hk3$ and $2kl$ planes ($14kl$ in the expanded cell), as shown in panels (c) and (f), respectively. The inset in (f) provides a magnified view at $l$ = 3, where the satellites along ky are clearly observed.
@@
We note that due to the anomalous intensity and satellite reflections occurring at $l$ = odd planes, a refinement using a structural model with a lattice parameter of $c/2$ relative to the HfFe6Ge6-type cell yields an improved wR of 10.05% (details in the Appendix D). This structural model corresponds to a YCo6Ge6-type cell, which is related to the HfFe6Ge6-type structure by a $c/2$ translation, a halving of the unit cell along $c$ and a resulting 50% occupancy at the U and Sn atomic positions along the $c$ axis (positions $1a$ and $2c$ in Table 3 of Appendix D, Fig. 10). The halving of the unit cell along $c$ means that the $l$ = odd planes are no longer integers, so that Bragg reflections are not expected there. This structure has been reported for the compound UFe6Ge6 [30], and was originally reported for DyFe6Sn6 and HoFe6Sn6 [31]. However, similar to the latter cases, we found that this YCo6Ge6-type model does not accurately represent the UV6Sn6 structure.
@@
We discovered that the satellite reflections observed in the diffraction patterns of UV6Sn6 can be accounted for by a supercell arising from a commensurate modulation of the HfFe6Ge6-type parent cell. The modulation corresponds to a periodic expansion of the parent unit cell by a factor of seven along the $a$ and $b$ axes (Fig. 1d-e), yielding a $7\times 7$ superstructure in the $ab$ plane. There is no evidence of cell expansion along $c$. The crystal symmetry remains hexagonal, $P6/mmm$, and the supercell lattice parameters are $a=b=38.548(2)$; $c=9.154(1)$, with $\alpha=\beta=90^{\circ};\gamma=120^{\circ}$. As detailed in the Appendix D, this modulation arises from the ordering of hexagonal cavities occupied by uranium in UV6Sn6 (Fig. 1c,d,e). This type of ordering is observed in other RT6X6 compounds, leading to patterns that extend the structures into supercells with modulations ranging from 2 to 68 parent unit cells [21, 22]. These atomic arrangements are explained by electrostatic and steric (size-related) effects, which impose preferences regarding the simultaneous occupation of neighboring cavities along $c$ [22].
@@
In the extended cell of UV6Sn6, the uranium-occupied hexagonal cavities shift along the $c$ axis by $c/2$ from their original position in the HfFe6Ge6-type parent structure, as illustrated in Fig. 1d. Labelling the shifts of the hexagonal cavities as $z=0$ for the original position and $z=$ for the $c/2$ shifted (in orange and blue respectively in Fig. 1d), the resulting pattern along the $a$ and $b$ crystallographic axes is 0, 0, ½, ½, ½, ½, 0. There is nearly an equal number of $z$ = 0 and $z$ = ½ sites in the supercell (25 and 24). This further explains why a refinement against the YCo6Ge6-type cell, with 50% occupancy at the U sites, yields a reasonable fit. Refinement of this supercell is challenging due to the large number of parameters and data points. However, the simulated diffraction patterns for our proposed UV6Sn6 expanded cell accurately replicate the diffracted spots observed in the reconstructed precession images, as demonstrated in Figs. 2c,f. Some of these superstructure reflections associated with the supercell observed in the neutron diffraction experiments are shown in Appendix C.

### v3

@@ Abstract
We report on the synthesis of single crystals of the kagome compound, UV6Sn6, and present the results of magnetization, electrical resistivity, heat capacity, x-ray, and neutron diffraction experiments to characterize the structure and magnetic properties. UV6Sn6 crystallizes in a large supercell of the HfFe6Ge6 parent structure with an hexagonal symmetry in which some of the U atoms are shifted by $c/2$ in an ordered fashion. Below $T_{N1}\approx 29$ K, an incommensurate magnetic structure with a temperature-dependent wave vector $(0,0,k_{z})$ is observed. Below $T_{N2}=23.5$ K, the wave vector locks in to $(0,0,0.5)$, forming an antiferromagnetic ground state. The U moments align along the $c$ axis retaining a large magnetic anisotropy. These findings highlight the role of the $5f$ orbitals from uranium in this structural family in driving both magnetic ordering and structural modulation and distinguish UV6Sn6 from its lanthanide-based analogs.
@@ I Introduction
Materials containing kagome networks have attracted a lot of attention for the study of topologically nontrivial phases, and intertwined orders [1, 2, 3]. Among them, the 166 family of compounds, characterized by the general formula $RT_{6}X_{6}$, represents a fertile platform for investigating the rich interplay of topology, electron correlations, and magnetism. The R element can be a rare-earth element or Li, Mg, Ca, Ti, Zr, Hf, T is a transition metal element, and X is typically a tetrel element such as Sn or Ge. In these materials, the kagome lattice, a two-dimensional network of corner-sharing triangles, facilitates the emergence of exotic phenomena, including charge density waves (CDWs) [4, 5, 6], phonon softening [7, 8, 9], gapped Dirac fermions and surface states [10, 11, 12], a nodal ring [13], anomalous and topological Hall effects [14, 15, 16], anomalous Nernst effect [17, 18], and an extraordinary structural diversity [19, 20, 21, 22, 23, 24, 25, 26].
@@ II Methods
Single crystals of UV6Sn6 were grown using a self-flux method similar to that reported for YV6Sn6 [11]. Natural uranium ingots (99.27% 238U, 0.72% 235U, trace 234U, from an internal stock at CEA Grenoble) were cut in $\sim$0.5 g pieces and cleaned by electrolytic etching to remove the oxides on the surface [27]. Some crystals were grown at UC Davis following the same procedure with natural uranium from the NBL Program office (99.24% 238U, 0.72% 235U, trace 234U). Vanadium pieces (99.9%, Aldrich) and tin shots (99.999%, Thermo Fisher Scientific) were cleaned with 30 vol.HCl (Sigma-Aldrich). The clean metals were introduced in a Canfield crucible set [28, 29] in the 1:6:20 molar ratio, and sealed in a silica ampoule with 1/3 atm argon gas. The ampoule was heated to 1150 °C in 7 h, held there for 36 h, slowly cooled to 1050 °C in 287 h. At 1050 °C, the formed crystals were decanted by removing the ampoule from the furnace, inverting it, and immediately spinning it in a centrifuge [30]. Only crystals of UV6Sn6 were obtained with this temperature profile. The experiments with cooling to 780 °C resulted in additional single crystals of USn3. The as-synthesized UV6Sn6 single crystals exhibit a metallic gray color and hexagonal pyramidal morphology with edge lengths of 0.3–0.8 mm (volume $\sim$0.02–0.33 mm3, Fig. 1d). Laue x-ray diffraction confirms that the $c$ axis is oriented perpendicular to the hexagonal base, while the $a$ and $b$ axes extend toward the in-plane hexagonal vertices.
@@
Single-crystal x-ray diffraction data were collected at 100–150 K using the four-circle Kappa geometry diffractometers: Rigaku XCallibur (Rigaku CrysAlis CCD software), and Bruker D8 Venture (APEX software). In both instruments, a Mo $\text{Mo }K_{\alpha}$ radiation source was used (0.71073 Å). The structural refinement was carried out using the crystallography software package GSAS-II [31].
@@
Magnetization measurements were performed using an MPMS3 SQUID magnetometer (Quantum Design) in VSM scan mode. The crystallographic axes of the sample were determined using Laue diffraction followed by polishing to obtain a flat surface perpendicular to the $c$ axis.
@@
Resistivity measurements were performed in a physical property measurement system (PPMS, Quantum Design) using a four point lock-in technique; 15 $\mu$m Au wires were spot welded to the sample for electrical contacts. Due to the irregular shape of the sample and the unperfect alignment of the electrical contacts, absolute values of the resistivity were not determined. The specific heat was measured using the PPMS with the heat pulse relaxation method; heat pulses of 0.5, 1, and 2% of the actual temperature gave similar results. The samples for specific heat had masses of 1.45 and 0.9 mg, thus the specific heat of the sample was small compared to the addenda contribution.
@@
Neutron diffraction experiments were carried out at Institut Laue Langevin, Grenoble on the two-axis diffractometer D23 and the large solid angle detector diffractometer XtremeD. A single-crystal sample with a mass of 1.48 mg was pre-aligned using x-ray diffraction (Laue method) with the $b$ axis oriented vertically, so the horizontal scattering plane was ($a^{*}$, $c^{*}$) in a hexagonal setting. D23 is equipped with a lifting detector reaching 30° and neutrons of wavelength 1.28 Å and 2.36 Å were used with a copper monochromator and a pyrolitic graphite (PG) monochromator, respectively. In this latter case, a PG filter was inserted in the incoming beam to remove higher order contamination. On XtremeD, the wavelength of 2.45 Å provided by a PG monochromator was used, the angular coverage of the detector was of 130° in the horizontal plane and 24° in the vertical plane. A PG filter was also installed. An oscillating radial collimator was placed in front of the detector bank to suppress the sample environment contribution to the collected data and reduce background. For both experiments, the sample was inserted in an helium flow cryostat. The MAG2POL software was used to refine the nuclear and magnetic structures [32].
@@ III.1 Crystal structure
Figure 1: The HfFe6Ge6-type parent structure of UV6Sn6, where the V atoms (red) form a kagome network (a), and the U atoms (blue) form a triangular lattice (b). The V kagome and U triangular lattices arrange in layers to form the structure depicted in (c), where the V and Sn atoms around U form a hexagonal cavity represented as a polyhedron. A picture of a representative single-crystal sample on a millimeter grid is shown in (d). The ordering of the cavities can lead to patterns such as the one shown in (e). These patterns extend the structures of $RT_{6}X_{6}$ compounds into supercells, like the one depicted in (f). The structure of UV6Sn6 can be described as a supercell where a periodic modulation of the HfFe6Ge6-type parent cell—marked in black in (f), yields a 7 × 7 superstructure on the $ab$ plane. The hexagonal cavities occupied by a U shift along the $c$ axis by $c/2$ yielding the pattern shown in (e). This pattern is identical along the $b$ crystallographic axis ($a=b$).
@@
Although the goodness-of-fit value of 1.76 (GOF, Table 1 in Appendix A) suggests reasonable agreement between the measured and calculated average intensities, the high value of the weighted profile factor, $wR2$ = 20.76%, reveals a significant discrepancy between the observed and calculated structure factors.
@@
Careful examination of both the x-ray and neutron diffraction data revealed that the discrepancies with the HfFe6Ge6-type model arise from very weak intensity at some families of Bragg reflections and the presence of additional satellite reflections around them. Notably, it was at the instrument XtremeD at ILL that we first observed these satellite reflections (see Appendix B), motivating further x-ray diffraction experiments. The satellite reflections are particularly evident in the reconstruction of reciprocal space (calculated precession images) centered on planes $hkl$, with $l$ = 3 and $h$ = 2, from x-ray diffraction data (Fig. 2b,e). On the $hk3$ plane (Fig. 2b), we clearly observe groups of satellite reflections around the expected Bragg reflections (e.g. 003, 103, and 013 marked in blue in Fig. 2a), as well as along the $h$ and $k$ directions, forming a triangular grid. Meanwhile, on the $2kl$ plane (Fig. 2e), the apparent discrepancy in the expected peak intensity and the presence of satellite reflections are evident for all odd values of $l$ (i.e., $l$ = 1, 3, 5, etc.).
@@
Figure 2: Comparison of the calculated and observed reconstructions of the reciprocal space centered on planes $hk3$ and $2kl$ for UV6Sn6. The calculated diffraction patterns in (a) and (d) show the expected Bragg spots (reflections, marked in blue) for a HfFe6Ge6-type cell on planes $hk3$ and $2kl$, respectively. The observed $hk3$ pattern in (b) shows undetectable intensity at the expected Bragg spots and the presence of six satellite reflections around them. Additional satellite reflections are observed along $h$ and $k$ directions. The observed $2kl$ pattern in (e), indexed in the HfFe6Ge6-type cell, shows discrepancies in the intensity of the expected Bragg reflections, as well as satellite peaks on either side of the Bragg spots along ky. These unexpected features are true for every $l$ = odd value (i.e., $l$ = 1, 3, 5, etc.). Our simulated diffraction patterns for an expanded $7\times 7$ cell (Fig. 1e-f) replicate the observed precession images for the $hk3$ and $2kl$ planes ($14kl$ in the expanded cell), as shown in (c) and (f), respectively. The inset in (f) provides a magnified view at $l$ = 3, where the satellites along ky are observed.
@@
We note that due to the anomalous intensity and satellite reflections occurring at $l$ = odd planes, a refinement using a structural model with a lattice parameter of $c/2$ relative to the HfFe6Ge6-type cell yields an improved $wR2$ of 10.05% (details in the Appendix C). This structural model corresponds to a YCo6Ge6-type cell, which is related to the HfFe6Ge6-type structure by a $c/2$ translation, a halving of the unit cell along $c$ and a resulting 50% occupancy at the U and Sn atomic positions along the $c$ axis (positions $1a$ and $2c$ in Table 3 of Appendix C, Fig. 10). This cell type is therefore often referred to as a disordered Y0.5Co3Ge3-type structure. During this study, a monoclinic distortion of the YCo6Ge6-type cell with an incommensurate structural wave vector was independently reported for UV6Sn6 with even better $wR2$ [33]. The halving of the unit cell along $c$ means that the $l$ = odd planes are no longer integers, so Bragg reflections are not expected there, requiring an additional incommensurate structural wave vector. The YCo6Ge6-type structure has also been reported for the compound UFe6Ge6 [34], and was originally reported for DyFe6Sn6 and HoFe6Sn6 [35]. However, similar to the latter cases, we found that this YCo6Ge6-type model does not accurately represent our UV6Sn6 structure.
@@
We discovered that the satellite reflections observed in the diffraction patterns of UV6Sn6 can be accounted for by a supercell arising from a modulation of the HfFe6Ge6-type parent cell. The modulation corresponds to a periodic expansion of the parent unit cell by a factor of 7 along the $a$ and $b$ axes (Fig. 1d-e), yielding a $7\times 7$ superstructure in the $ab$ plane. There is no evidence of cell expansion along $c$. The crystal symmetry remains hexagonal, $P6/mmm$, and the supercell lattice parameters are $a=b=38.548(2)$; $c=9.154(1)$, with $\alpha=\beta=90^{\circ};\gamma=120^{\circ}$. As detailed in the Appendix D, this modulation arises from the ordering of hexagonal cavities occupied by uranium in UV6Sn6 (Fig. 1c,e,f; these cavities are also referred to as voids in the literature). This type of ordering is observed in other RT6X6 compounds, leading to patterns that extend the structures into supercells with modulations ranging from 2 to 68 parent unit cells [21, 22]. These atomic arrangements are explained by electrostatic and steric (size-related) effects, which impose preferences regarding the simultaneous occupation of neighboring cavities along $c$ [22]. We note that in contrast to other modulated structures such as $R$Fe6Sn6 with $R$ = Er, Y, Tb [22], the modulation in our UV6Sn6 superstructure retains the hexagonal symmetry. An orthorhombic or monoclinic cell with a single $q$ modulation such as the one proposed in Ref. [33] only accounts for a pair of satellite reflections rather than the six we observe in the $hk3$ plane in Fig. 2b. It is possible that our observations result from triply twinned crystals with the monoclinic structure proposed by Thomas et al [33]. From diffraction data collected from various single crystals of different sizes–including some of dimensions comparable to those reported for x-ray diffraction in Ref. [33], we consistently observe satellite reflections that retain the global hexagonal symmetry within the $ab$ plane. This is also observed in the neutron diffraction data shown in Fig. 9 of Appendix B. It is thus likely that we have crystallized a high-temperature structure with a hexagonal symmetry, different from the monoclinic structure reported in Ref. [33]. This is possible considering the different temperature profile compared to that in Ref. [33]. A sensitivity to the thermal treatment for the stabilization of different crystal structures has been reported for other $RT_{6}X_{6}$ phases [36, 37], and may be critical for understanding the tunability of the magnetic and charge ordered phases of $R$Mn6Sn6, $R$Fe6Ge6 and $Ln$Nb6Sn6 ($Ln$ = lanthanide), respectively [3, 38, 26].
@@
In our proposed extended cell of UV6Sn6, the uranium-occupied hexagonal cavities shift along the $c$ axis by $c/2$ from their original position in the HfFe6Ge6-type parent structure, as illustrated in Fig. 1e. Labelling the shifts of the hexagonal cavities as $z=0$ for the original position and $z=$ for the $c/2$ shifted (in orange and blue, respectively, in Fig. 1e), the resulting pattern along the $a$ and $b$ crystallographic axes is 0, 0, ½, ½, ½, ½, 0. There is nearly an equal number of $z$ = 0 and $z$ = ½ sites in the supercell (25 and 24). This further explains why a refinement against the YCo6Ge6-type cell, with 50% occupancy at the U sites, yields a reasonable fit. Refinement of this supercell is challenging due to the large number of parameters and data points. However, the simulated diffraction patterns for our proposed UV6Sn6 expanded cell accurately replicate the diffracted spots observed in the reconstructed precession images, as demonstrated in Figs. 2c,f.
As noted above, the superstructure of UV6Sn6 is significantly different from other $RT_{6}X_{6}$ compounds. While the atomic sizes and lattice parameters in UV6Sn6 are similar to those of other $R$V6Sn6 compounds (where $R$ = Y, Sm, Gd, Dy-Lu), most of these compounds typically adopt either the undistorted, non modulated HfFe6Ge6-type structure or the partially disordered SmMn6Ge6-type structure (with the YCo6Ge6-type at the fully disordered extreme) [37, 26]. We observe that, generally, larger radii at $R$ seem to lead to modulated structures, which could be explained from the expansion of the hexagonal cages (Fig. 1c), which imposes additional steric constraints. The superstructure in UV6Sn6 might also be driven by the spatially extended nature of the U $5f$ orbitals.
As will be discussed below, we do not find evidence of a CDW state in UV6Sn6. This contrasts with the closely related compound ScV6Sn6 [4, 6]. The radius of Sc, $\sim 0.89$, is considerably smaller than that of U, $\sim 1.14$, (VIII-coordinate Shannon radius [39] as compared in Ref. [26]). An additional consequence of the expansion of the hexagonal cavities when occupied by atoms with increasing radii is the reduction of Sn–Sn distances. This bond shortening limits the displacement of Sn along the $c$ axis, thus suppressing the “rattling” $R$-Sn-Sn-$R$ bond modulation, identified as a potential CDW instability in ScV6Sn6 [6, 26, 40].